Find the center of mass of a triangular sheet whose vertices are (8, 0), (2, 6) and (2, 0), whose density function is

p (x, y) = x + y + 1







1
Expert's answer
2021-09-28T15:52:59-0400

Vertices:

"(0,8),(2.6),(2,0)"


Graphs:

"(0,8),(2.6)" :

"8=b"

"6=2a+b\implies a=-1"

"y=8-x"


"(2.6),(2,0)"

"x=2"


"(0,8),(2,0)"

"b=8"

"0=2a+b\implies a=-4"

"y=8-4x"



mass:

"m=\iint \rho(x,y)dA=\int^{x=2}_{x=0}\int^{y=8-x}_{y=8-4x}(x+y+1)dydx="


"=\int^{x=2}_{x=0}(xy+y^2/2+y)|^{y=8-x}_{y=8-4x}dx="


"=\int^{x=2}_{x=0}(x(8-x)+(8-x)^2/2+8-x-x(8-4x)-(8-4x)^2/2-8+4x)dx="


"=\int^{x=2}_{x=0}(-4.5x^2+27x)dx=(27x^2/2-1.5x^3)|^2_0=42"


"M_x=\iint y\rho(x.y)dA=\int^{x=2}_{x=0}\int^{y=8-x}_{y=8-4x}y(x+y+1)dydx="


"=\int^{x=2}_{x=0}(y^2x/2+y^3/3+y^2/2)|^{y=8-x}_{y=8-4x}dx="


"=\int^{x=2}_{x=0}(x(8-x)^2/2+(8-x)^3/3+(8-x)^2/2-x(8-4x)^2/2-(8-4x)^3/3-"


"-(8-4x)^2/2)dx=\int^{x=2}_{x=0}(32x-8x^2+x^3/2+512/3-64x+8x^2-x^3/3+"


"+32-8x+x^2\/2-32x+32x^2-8x^3-512\/3+256x-128x^2+64x^3\/3-"


"-32+32x-8x^2)dx=\int^{x=2}_{x=0}(216x-103.5x^2+13.5x^3)dx="


"=(108x^2-34.5x^3+3.375x^4)|^2_0=210"


"M_y=\iint x\rho(x.y)dA=\int^{x=2}_{x=0}\int^{y=8-x}_{y=8-4x}x(x+y+1)dydx="


"=\int^{x=2}_{x=0}(y^2x/2+x^2y+xy)|^{y=8-x}_{y=8-4x}dx="


"=\int^{x=2}_{x=0}((8-x)^2x/2+x^2(8-x)+x(8-x)-(8-4x)^2x/2-"


"-x^2(8-4x)-x(8-4x))dx=\int^{x=2}_{x=0}(32x-8x^2+x^3/2+8x^2-x^3+"


"+8x-x^2-32x+32x^2-8x^3-8x^2+32x^3-8x+4x^2)dx="


"=(9x^3+5.75x^4)|^2_0=164"


The center of mass:


"\overline{x}=\frac{M_y}{m}=\frac{164}{42}=\frac{82}{21}=3.90"


"\overline{y}=\frac{M_x}{m}=\frac{210}{42}=5.00"